# CURATRON: Complete Robust Preference Data for Robust Alignment of Large Language Models

### v1

@@ 8 Conclusion
Our study examines how missing information and distorted feedback can impact LLMs, potentially compromising their performance in terms of alignment with human values. We have proposed a robust algorithm for provably correct and efficient ranking responses in the BTL, LR, and general binary choice models. This robust ranking data is then input in the PL step. Further, we also handled the partially observed setting, wherein only some response pairs are compared, by integrating matrix completion techniques into our robust learning algorithm. In all cases, we provided statistical and computational guarantees using novel techniques. Through our comprehensive analysis, we hope to contribute to the ongoing discussion on AI safety by helping to create and scale LLMs/AGI models that align with human values and expectations. Some future research directions include tightening the recovery results for partially observed settings under weaker conditions (possibly using noisy-case extensions of (Yi et al.,, 2016)), exploring other notions of adversarial noise, and understanding the minimax optimal rates for ranking estimators under various noise models. We also plan to study the parametric non-active pairwise ranking setting, studying lower bounds and practical algorithms in the active setting similar to (Heckel et al.,, 2016). Furthermore, it would be interesting to investigate whether we can extend this approach to solve the entity corruption problem in retrieval models, as shown in (Naresh et al.,, 2022). Another research direction could be defining an alignment framework that expands DPO to various objective functions based on Rank Centrality (Negahban et al.,, 2017). Finally, we aim to examine the relationship between robust PL and model capacity, as this can shed light on the trade-offs between model complexity and generalization performance.
Appendix A Prompt template and answer of LLM judge
We evaluate various responses from different Language Models (LLMs) using GPT-4-1106-preview (OpenAI et al.,, 2023) as the judge. We have adapted the prompt template from MT-Bench (Zheng et al.,, 2023), but instead of scoring each candidate separately, we score all models in a single turn. Please note that the results generated by GPT-4-1106-preview judge may not be the same when the model gets updated. To ensure that the judge’s response is deterministic, we have set the temperature parameter to 0.
System prompt: You are a helpful assistant. Prompt Template: [Instruction] Please act as an impartial judge and evaluate the qualities of the responses provided by many AI assistants to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Begin your evaluation by providing a short explanation to each of the assistant responses. Be as objective as possible. Your judgment should be based solely on the content of the answers, disregarding factors like the position of the responses, the length of the responses, or the names of the assistants. After providing your explanation, you must rate the responses on a scale of 0 to 1 by strictly following this format: Final verdict for each assistant (full list): {‘Assistant A’: $score_{a}$, ’Assistant n’: $score_{n}$} For example: Final verdict for each assistant (full list): {‘Assistant A’: 0.30, ‘Assistant B’: 0.19, ‘Assistant C’: 0.75} [The Start of Assistant A’s Answer] Assistant A’s Answer [The End of A’s Answer] [The Start of Assistant n’s Answer] Assistant n’s Answer [The End of n’s Answer] Judge’s Answer: Final verdict for each assistant (full list):
{‘Assistant A’: 0.90, ‘Assistant B’: 0.95, ‘Assistant C’: 0.80, ‘Assistant D’: 0.75, ‘Assistant E’: 0.85, ‘Assistant F’: 0.70, ‘Assistant G’: 0.60, ‘Assistant H’: 0.65, ‘Assistant I’: 0.55, ‘Assistant J’: 0.40, ‘Assistant K’: 0.45, ‘Assistant L’: 0.50, ‘Assistant M’: 0.20, ‘Assistant N’: 0.35, ‘Assistant O’: 0.30}
Explanation of Scores: A scored well as it provides a structured, detailed, and vivid account of the Hawaiian experience, including cultural and natural attractions. B earned the highest score for a well-written, engaging narrative combining in-depth cultural experiences with comprehensive overviews of natural attractions, presenting it all in an engaging manner. C’s response was immersive, providing a daily itinerary format which gives readers a chronological perspective but loses some detail and depth in the descriptions, resulting in a slightly lower score. D’s response was engaging, offering personal reflections and detailed descriptions of the experiences, but it could capitalize on a more informative angle regarding cultural aspects. E crafted an elaborate and expressive narrative with a strong focus on cultural elements and natural beauty. It’s highly descriptive, which enhances its appeal. F offered a nice overview with personal insights but was more generic compared to others, with less distinctive detail on the cultural elements and attractions. G provided a good variety of information but lacked a personal touch and the engaging storytelling present in others, hence a middle-of-the-road score. H put forth a decent attempt to cover cultural experiences and attractions, but the content was less rounded and not as rich in personal perspective. I’s storytelling lacked in-depth detail and read more like an overview, resulting in a lower rating compared to others that provided richer narratives. J’s response was brief and lacked the depth, detail, and storytelling elements found in the higher-scoring responses. K focused more on personal experiences rather than providing informative cultural insights, which resulted in a score below others that balanced both elements better. L offered an informative narrative but was less dynamic and less vivid in description than higher-scoring answers. M’s response was too factual and lacked the storytelling and descriptive engagement present in other responses, leading to a lower score. N provided a decent overview but lacked the richness of storytelling and description, leaving it towards the lower end of the scale. O’s draft was informative but lacked the detail and narrative quality that made other responses stand out, resulting in a lower score.

### v2

@@ 8 Discussion
Some other future research directions include implementing our algorithms tensor-wise for batch processing with GPU to improve speed, tightening the recovery results for partially observed settings under weaker conditions (possibly using noisy-case extensions of (Yi et al.,, 2016)), exploring other notions of adversarial noise, and understanding the minimax optimal rates for ranking estimators under various noise models. We also plan to study the parametric non-active pairwise ranking setting, studying lower bounds and practical algorithms in the active setting similar to (Heckel et al.,, 2016). Furthermore, it would be interesting to investigate whether we can extend this approach to solve the entity corruption problem in retrieval models, as shown in (Naresh et al.,, 2022). Another research direction could be defining an alignment framework that expands DPO to various objective functions based on Rank Centrality (Negahban et al.,, 2017). Finally, we aim to examine the relationship between robust PL and model capacity, as this can shed light on the trade-offs between model complexity and generalization performance.
9 Conclusions
Our study examines how missing information and distorted feedback can impact LLMs, potentially compromising their performance in terms of alignment with human values. We have proposed rigorous algorithms for provably correct and efficient ranking responses in the BTL, LR, and general binary choice models. In all cases, we provided statistical and computational guarantees using novel techniques. Additionally, we propose a novel data augmentation technique to improve the performance of our algorithms further. We aim to contribute to the ongoing discussion on AI safety/Responsible AI by helping in the development and scaling of LLMs/AGI models that align with human values and expectations.